OpenStudy (anonymous):

Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (luigi0210):

Part 1: y+2 = 2(x-3) Part 2: y+2 = 2x-6 y=2x-8 Part 3: -2x+y=-8 2x-y = 8

OpenStudy (anonymous):

Thank you